What is the wavelength of electromagnetic radiation which has a frequency of 7.215 x 10-6 s? (c = 2.998 x 108 m/s)?
Answer #1

Frequency t wavelength are related is, (2) Frequency = speed of light (c). Frequency of Wavelength (2)2= 7.2158 10661 c= 2.998% 108 m/s x = 2.998 x 10 m/s 7.215 x 1016 st 1 = 2.998 * 108-16m 7.215 Wavelength = 0-416 x 10-8m 4.1